# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ 1.3. Organization of the paper
In Section 2, we introduce the background tools necessary for the rest of the paper. Specifically, §2.1 provides a short introduction to the geometry and Fourier analysis of the unit tangent bundle. It also introduces the new objects needed to deal with the divergence of the generalized thermostats. In §2.2, we review the complex geometry and harmonic analysis on a surface, while §2.3 delves into hyperbolic dynamics and tensor tomography.
In Section 3, we explain how a smooth orbit equivalence acts on holomorphic differentials, and we establish the pairing formula needed to show that period matrices are preserved. We then present the proofs of our main results in Section 4.
Appendix A delves into the question of finding distributional solutions, with prescribed Fourier modes, of the relevant transport equation for a generalized thermostat.
@@ Acknowledgements
I would like to thank my advisor, Gabriel Paternain, for suggesting this project and guiding me while working on it.
@@ 2. Preliminaries
In what follows, $(M,g)$ is a smooth closed oriented Riemannian surface, and we take an arbitrary $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$. Whenever we use additional assumptions, it will be clearly stated in the result statements. We will sometimes need a second generalized thermostat $(M,\tilde{g},\tilde{\lambda})$. All the objects depending on the metric will then be labeled accordingly. Finally, we denote by $\phi:\tilde{S}M\to SM$ a smooth orbit equivalence between the thermostats $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$. Once again, we will specify when we assume it to be isotopic to the identity.
@@ 2.1. Unit tangent bundle of the surface
We review some basics of the unit tangent bundle $\pi:SM\to M$ defined by
$SM:=\{(x,v)\in TM\mid\|v\|_{g}=1\}.$
@@ 2.1.1. Geometry of $SM$
As previously, let $X$ be the geodesic vector field on $SM$, and let $V$ be the vertical vector field generating the circle action on the fibers. We define $H:=[V,X]$. The vector fields $\{X,H,V\}$ form an orthonormal basis on $SM$ for the Sasaki metric (the natural lift of $g$ to $SM$). We set $\mathbb{H}:=\mathbb{R}H$ and $\mathbb{V}:=\mathbb{R}V$. We also note that the geodesic vector field splits into $X=\eta_{+}+\eta_{-}$ where $\eta_{\pm}$ are the raising and lowering Guillemin-Kazhdan operators given by
$\eta_{\pm}:=\dfrac{1}{2}(X\mp iH).$ (2.1)
The Liouville $1$-form $\alpha\in\mathcal{C}^{\infty}(SM,T^{*}(SM))$ is defined by $\alpha(X)=1$ and $\alpha(H)=\alpha(V)=0$. It is invariant by the geodesic flow in the sense that $\mathcal{L}_{X}\alpha=0$. The $2$-form $d\alpha$ is non-degenerate on the contact plane $\mathbb{H}\oplus\mathbb{V}$, and it satisfies $\iota_{X}d\alpha=0$. Hence
@@
is a volume form invariant by the geodesic flow. We call it the Liouville volume form. It corresponds to the Riemannian volume form induced by the Sasaki metric on $SM$. From now on, the $L^{2}$ space on $SM$ is defined as $L^{2}(SM):=L^{2}(SM,\mu)$.
We also define the $1$-forms $\beta,\psi$ on $SM$ by $\beta(H)=1=\psi(V)$ and $\beta(X)=\beta(V)=0=\psi(X)=\psi(H)$. It is easy to check that $d\alpha=\psi\wedge\beta$ so that $\mu=\alpha\wedge\beta\wedge\psi$. We set $(\mathbb{R}X)^{\ast}:=\mathbb{R}\alpha$, $(\mathbb{R}F)^{\ast}:=\mathbb{R}\lambda$, $\mathbb{H}^{\ast}:=\mathbb{R}\beta$, and $\mathbb{V}^{\ast}:=\mathbb{R}\psi$. We refer to [PSU23, Chapter 3] for further details on the geometric structure on $SM$.
@@ 2.1.2. Appearance of the divergence
The key difference between generalized thermostats and geodesic or magnetic flows is that the generating vector field $F$ might not preserve the Liouville volume form $\mu$. Recall that the divergence of the vector field $F$ with respect to the volume form $\mu$ is the function $\mathrm{div}_{\mu}F\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ uniquely defined by
@@
Let $(M,g,\lambda)$ be a generalized thermostat. Then, we have:
@@
In the geodesic and magnetic cases, we have $V(\lambda)=0$, so the Liouville volume form is preserved. Another way in which the divergence manifests itself is when calculating the adjoint operators with respect to the $L^{2}$ inner product on $SM$:
$F^{*}=-(F+V(\lambda)),\qquad H^{*}=-H,\qquad V^{*}=-V.$
This is relevant when extending differential operators to act on the space of distributions. Recall that any differential operator $P$ with smooth real-valued coefficients acts on a distribution $u\in\mathcal{D}^{\prime}(SM)$ by duality, that is $\langle Pu,\varphi\rangle_{\mathcal{D}^{\prime}(SM)}:=\langle u,P^{*}\varphi\rangle_{\mathcal{D}^{\prime}(SM)}$ for any $\varphi\in\mathcal{C}^{\infty}(SM)$. The subspace of distributional solutions to the transport equation
$\mathcal{D}^{\prime}_{\text{tr}}(SM):=\{u\in\mathcal{D}^{\prime}(SM)\mid(F+V(\lambda))u=0\}$
thus corresponds to the distributions $u\in\mathcal{D}^{\prime}(SM)$ such that $\langle u,F\varphi\rangle_{\mathcal{D}^{\prime}(SM)}=0$ for all $\varphi\in\mathcal{C}^{\infty}(SM)$. If $V(\lambda)=0$, these are simply the distributions invariant by the flow.
@@ 2.1.3. Divergence and smooth orbit equivalences
It will prove important to understand how the divergence of a system interacts with smooth orbit equivalences.
@@
Let $N$ and $\tilde{N}$ be two orientable manifolds endowed with nowhere-vanishing volume forms $\mu$ and $\tilde{\mu}$, and smooth vector fields $Y$ and $\tilde{Y}$. Suppose $\phi:\tilde{N}\to N$ is a smooth orbit equivalence between the flows generated by $\tilde{Y}$ and $Y$. If we write $\phi_{*}\tilde{Y}=cY$ with $c\in\mathcal{C}^{\infty}(N,\mathbb{R}_{>0})$ and $\phi^{*}\mu=(\det\phi)\tilde{\mu}$ with $\det\phi\in\mathcal{C}^{\infty}(\tilde{N},\mathbb{R})$, then
$(\det\phi)\phi^{*}(\mathrm{div}_{\mu}Y)=\tilde{Y}\left(\dfrac{\det\phi}{\phi^{*}c}\right)+\dfrac{\det\phi}{\phi^{*}c}\mathrm{div}_{\tilde{\mu}}\tilde{Y}.$
In particular, if $\phi$ preserves the orientation, i.e., $\det\phi>0$, then
$(\phi^{*}c)\phi^{*}(\mathrm{div}_{\mu}Y)=\tilde{Y}\left(\ln\left(\dfrac{\det\phi}{\phi^{*}c}\right)\right)+\mathrm{div}_{\tilde{\mu}}\tilde{Y}.$
@@
$=d\left(\dfrac{\det\phi}{\phi^{*}c}\iota_{\tilde{Y}}\tilde{\mu}\right)$
$=d\left(\dfrac{\det\phi}{\phi^{*}c}\right)\wedge\iota_{\tilde{Y}}\tilde{\mu}+\dfrac{\det\phi}{\phi^{*}c}d(\iota_{\tilde{Y}}\tilde{\mu})$
$=\left(\tilde{Y}\left(\dfrac{\det\phi}{\phi^{*}c}\right)+\dfrac{\det\phi}{\phi^{*}c}\mathrm{div}_{\tilde{\mu}}\tilde{Y}\right)\tilde{\mu}.$
On the other hand, we also have
@@
In the geodesic and magnetic cases, the pullback $\phi^{*}$ of the smooth orbit equivalence $\phi:\tilde{S}M\to SM$ sends the space $\mathcal{D}^{\prime}_{\text{tr}}(SM)$ to $\mathcal{D}^{\prime}_{\text{tr}}(\tilde{S}M)$. More generally, however, the divergence term $V(\lambda)$ appearing in the transport equation breaks this down.
@@
$\mathcal{F}(SM):=\{\sigma\in\mathcal{D}^{\prime}(SM,\Lambda^{2}T^{*}(SM))\mid\iota_{F}\sigma=d\sigma=0\}.$
This set only depends on the foliation corresponding to $F$, i.e., it is invariant under time-changes, so we get a $\mathbb{C}$-linear isomorphism $\phi^{*}:\mathcal{F}(SM)\to\mathcal{F}(\tilde{S}M)$. The 2-form
$\omega:=\iota_{F}\mu$ (2.2)
@@
$=du\wedge\omega+u\mathcal{L}_{F}\mu$
$=du\wedge\omega+uV(\lambda)\mu$
@@
Therefore, $u\omega$ is closed if and only if $(F+V(\lambda))u=0$. Since $F$ never vanishes, any $2$-current $\sigma$ on $SM$ satisfying $\iota_{F}\sigma=0$ must be of the form $\sigma=u\omega$ for some $u\in\mathcal{D}^{\prime}(SM)$. ∎
Thanks to this identification, we can now define a map $\Phi:\mathcal{D}^{\prime}_{\text{tr}}(SM)\to\mathcal{D}^{\prime}_{\text{tr}}(\tilde{S}M)$ associated to the smooth orbit equivalence $\phi:\tilde{S}M\to SM$ via the following diagram:
@@
This point of view does not affect the wavefront set analysis.
@@
Let $q\in\mathcal{C}^{\infty}(\tilde{S}M,\mathbb{R}_{>0})$ be the function such that $\phi^{*}\omega=q\tilde{\omega}$. Then, we get
$\Phi u$ $=\tilde{L}^{-1}\phi^{*}Lu$
$=\tilde{L}^{-1}\phi^{*}(u\omega)$
$=\tilde{L}^{-1}(q\phi^{*}u\tilde{w})$
@@
Since multiplication by the nowhere-vanishing function $q$ is elliptic, we get the result by elliptic regularity (see [Hö03, Theorem 8.3.2]).
By the properties of wavefront sets under pullback operators (see [Hö03, Theorem 8.2.4] for instance), we thus obtain
$\textup{WF}(\Phi u)=d\phi^{T}(\textup{WF}(u))$
for all $u\in\mathcal{D}^{\prime}_{\textup{tr}}(SM)$, where $d\phi^{T}:T^{*}(SM)\to T^{*}(\tilde{S}M)$ is the symplectic lift of $\phi^{-1}$ to the cotangent bundles given by
$d\phi^{T}(y,\eta):=\left(\phi^{-1}(y),d\phi^{T}_{\phi^{-1}(y)}\eta\right),\quad(y,\eta)\in T^{*}(SM).$
2.1.4. Fourier decomposition
@@
$\mathcal{C}^{\infty}(SM)=\oplus_{k\in\mathbb{Z}}\Omega_{k},\qquad\Omega_{k}:=\{u\in\mathcal{C}^{\infty}(SM)\mid\,Vu=iku\}.$
This decomposition is orthogonal with respect to the $L^{2}$ inner product on $SM$ and with $\mathcal{C}^{\infty}$ being replaced by $L^{2}$. For any $u\in\mathcal{C}^{\infty}(SM)$, we shall write $u=\sum_{k\in\mathbb{Z}}u_{k}$, where each $u_{k}\in\Omega_{k}$ is given by
@@
$\langle u_{k},\varphi\rangle:=\langle u,\varphi_{-k}\rangle_{\mathcal{D}^{\prime}(SM)},\quad\varphi\in\mathcal{C}^{\infty}(SM),$
@@
It also worth noting that the ladder operators $\eta_{\pm}$ in (2.1) take their name from the fact that they act as raising/lowering operators on the Fourier decomposition, that is,

### v2

@@ 1.3. Organization of the paper
In §2, we introduce the background tools necessary for the rest of the paper. Specifically, §2.1 provides a short introduction to the geometry and vertical Fourier analysis of the unit tangent bundle. It also introduces the new objects needed to deal with the divergence of the thermostats. In §2.2, we review the complex geometry and harmonic analysis on a surface, while §2.3 delves into hyperbolic dynamics and tensor tomography.
In §3, we explain how a smooth orbit equivalence acts on holomorphic differentials and we establish the pairing formula needed to show that period matrices are preserved. We then present the proofs of our main results in §4.
Appendix A delves into the question of finding distributional solutions, with prescribed Fourier modes, of the relevant transport equation for a thermostat.
@@ Acknowledgements
I would like to thank my supervisor, Gabriel Paternain, for suggesting this project and guiding me while working on it. This research was supported by a Harding Distinguished Postgraduate Scholarship.
@@ 2. Preliminaries
In what follows, $(M,g)$ is a closed oriented Riemannian surface and we take an arbitrary $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$. Whenever we use additional assumptions, it will be clearly stated in the result statements. We will sometimes need a second thermostat $(M,\tilde{g},\tilde{\lambda})$. All the objects depending on the metric will then be labelled accordingly. Finally, we denote by $\phi:\widetilde{S}M\to SM$ a smooth orbit equivalence between the thermostats $(M,\tilde{g},\tilde{\lambda})$ and $(M,g,\lambda)$. Once again, we will specify when we assume it to be isotopic to the identity.
@@ 2.1. Unit tangent bundle of the surface
We review some basics concerning the unit tangent bundle of $M$, that is, the $3$-dimensional manifold defined as
$SM\coloneqq\left\{(x,v)\in TM\mid\|v\|_{g}=1\right\},$
together with its natural projection $\pi:SM\to M$.
@@ 2.1.1. Geometry of $SM$
As previously, let $X$ be the geodesic vector field on $SM$ and let $V$ be the vertical vector field generating the circle action on the fibres. We define $H\coloneqq[V,X]$. Then, $(X,H,V)$ is a positively oriented global frame for $T(SM)$, orthonormal with respect to the Sasaki metric (the natural lift of $g$ to $SM$). We set $\mathbb{H}\coloneqq\mathbb{R}H$ and $\mathbb{V}\coloneqq\mathbb{R}V$. We also note that the geodesic vector field splits into $X=\eta_{+}+\eta_{-}$, where $\eta_{\pm}$ are the raising and lowering Guillemin–Kazhdan operators given by
$\eta_{\pm}\coloneqq\dfrac{1}{2}(X\mp iH).$ (2.1)
The Liouville $1$-form $\alpha\in\mathcal{C}^{\infty}(SM,T^{*}(SM))$ is defined by the relations $\alpha(X)=1$ and $\alpha(H)=\alpha(V)=0$. It is invariant by the geodesic flow in the sense that $\mathcal{L}_{X}\alpha=0$. The $2$-form $d\alpha$ is non-degenerate on the contact plane $\mathbb{H}\oplus\mathbb{V}$ and it satisfies $\iota_{X}d\alpha=0$. Hence,
@@
is a nowhere-vanishing volume form invariant by the geodesic flow. We call it the Liouville volume form (and also use $\mu$ to denote the induced Liouville measure on $SM$). It coincides with the Riemannian volume form induced by the Sasaki metric on $SM$. From now on, the $L^{2}$ space on $SM$ is defined as $L^{2}(SM)\coloneqq L^{2}(SM,\mu)$.
We also define the $1$-forms $\beta,\psi$ on $SM$ by the relations $\beta(H)=1=\psi(V)$ and $\beta(X)=\beta(V)=0=\psi(X)=\psi(H)$. It is easy to check that $d\alpha=\psi\wedge\beta$ so that $\mu=\alpha\wedge\beta\wedge\psi$. We set $(\mathbb{R}X)^{\ast}\coloneqq\mathbb{R}\alpha$, $\mathbb{H}^{\ast}\coloneqq\mathbb{R}\beta$ and $\mathbb{V}^{\ast}\coloneqq\mathbb{R}\psi$. We refer to [PSU23, §3.5] for further details on the geometric structure of $SM$.
@@ 2.1.2. Appearance of the divergence
The key difference between thermostats and geodesic or magnetic flows is that the generating vector field $F$ might not preserve the Liouville volume form $\mu$. Recall that the divergence of the vector field $F$ with respect to the volume form $\mu$ is the function $\mathrm{div}_{\mu}F\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ uniquely characterized by
@@
Let $(M,g,\lambda)$ be a thermostat. Then, we have
@@
In the geodesic and magnetic cases, we have $V\lambda=0$, so the Liouville volume form is preserved. Another way in which the divergence manifests itself is when calculating the adjoint operators with respect to the $L^{2}$ inner product on $SM$:
$F^{*}=-(F+V\lambda),\qquad H^{*}=-H,\qquad V^{*}=-V.$
This is relevant when extending differential operators to act on the space of distributions. Recall that any differential operator $P$ with smooth real-valued coefficients acts on a distribution $u\in\mathcal{D}^{\prime}(SM)$ by duality, that is, $\langle Pu,\varphi\rangle_{\mathcal{D}^{\prime}(SM)}\coloneqq\langle u,P^{*}\varphi\rangle_{\mathcal{D}^{\prime}(SM)}$ for any $\varphi\in\mathcal{C}^{\infty}(SM)$. The subspace of distributional solutions to the transport equation
$\mathcal{D}^{\prime}_{\text{tr}}(SM)\coloneqq\left\{u\in\mathcal{D}^{\prime}(SM)\mid(F+V\lambda)u=0\right\}$
thus corresponds to the distributions $u\in\mathcal{D}^{\prime}(SM)$ such that $\langle u,F\varphi\rangle_{\mathcal{D}^{\prime}(SM)}=0$ for all $\varphi\in\mathcal{C}^{\infty}(SM)$. If $V\lambda=0$, these are simply the distributions invariant by the flow.
@@ 2.1.3. Divergence and smooth orbit equivalences
It is crucial to understand how the divergence of a system interacts with smooth orbit equivalences.
@@
Let $N$ and $\widetilde{N}$ be two oriented manifolds endowed with nowhere-vanishing volume forms $\mu$ and $\tilde{\mu}$, and smooth nowhere-vanishing vector fields $Y$ and $\widetilde{Y}$. Suppose that $\phi:\widetilde{N}\to N$ is a smooth orbit equivalence between the flows generated by $\widetilde{Y}$ and $Y$. If we write $\phi_{*}\widetilde{Y}=cY$ with $c\in\mathcal{C}^{\infty}(N,\mathbb{R}_{>0})$ and $\phi^{*}\mu=(\det\phi)\tilde{\mu}$ with $\det\phi\in\mathcal{C}^{\infty}(\widetilde{N},\mathbb{R})$, then
$(\det\phi)\phi^{*}(\mathrm{div}_{\mu}Y)=\widetilde{Y}\left(\dfrac{\det\phi}{\phi^{*}c}\right)+\dfrac{\det\phi}{\phi^{*}c}\mathrm{div}_{\tilde{\mu}}\widetilde{Y}.$
In particular, if $\phi$ preserves the orientation, that is, $\det\phi>0$, then
$\phi^{*}(c)\phi^{*}(\mathrm{div}_{\mu}Y)=\widetilde{Y}\left(\ln\left(\dfrac{\det\phi}{\phi^{*}c}\right)\right)+\mathrm{div}_{\tilde{\mu}}\widetilde{Y}.$
@@
$=d\left(\dfrac{\det\phi}{\phi^{*}c}\iota_{\widetilde{Y}}\tilde{\mu}\right)$
$=d\left(\dfrac{\det\phi}{\phi^{*}c}\right)\wedge\iota_{\widetilde{Y}}\tilde{\mu}+\dfrac{\det\phi}{\phi^{*}c}d(\iota_{\widetilde{Y}}\tilde{\mu})$
$=\left(\widetilde{Y}\left(\dfrac{\det\phi}{\phi^{*}c}\right)+\dfrac{\det\phi}{\phi^{*}c}\mathrm{div}_{\tilde{\mu}}\widetilde{Y}\right)\tilde{\mu}.$
However, we also have
@@
In the geodesic and magnetic cases, the pullback $\phi^{*}$ of the smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ sends the space $\mathcal{D}^{\prime}_{\text{tr}}(SM)$ to $\mathcal{D}^{\prime}_{\text{tr}}(\widetilde{S}M)$. More generally, however, the divergence term $V\lambda$ appearing in the transport equation breaks this down.
@@
$\mathcal{F}(SM)\coloneqq\left\{\sigma\in\mathcal{D}^{\prime}(SM,\Lambda^{2}T^{*}(SM))\mid\iota_{F}\sigma=d\sigma=0\right\}.$
This set only depends on the foliation corresponding to $F$, that is, it is invariant under time-changes, so we get a $\mathbb{C}$-linear isomorphism $\phi^{*}:\mathcal{F}(SM)\to\mathcal{F}(\widetilde{S}M)$. The 2-form
$\omega\coloneqq\iota_{F}\mu$ (2.2)
@@
$=du\wedge(\iota_{F}\mu)+u\mathcal{L}_{F}\mu$
$=\iota_{F}(du)\mu+uV(\lambda)\mu$
@@
Therefore, the $2$-current $u\omega$ is closed if and only if $(F+V\lambda)u=0$. Since $F$ is nowhere-vanishing, any $2$-current $\sigma$ on $SM$ satisfying $\iota_{F}\sigma=0$ must be of the form $\sigma=u\omega$ for some distribution $u\in\mathcal{D}^{\prime}(SM)$. ∎
Thanks to this identification, we can now define a map $\Phi:\mathcal{D}^{\prime}_{\text{tr}}(SM)\to\mathcal{D}^{\prime}_{\text{tr}}(\widetilde{S}M)$ associated to the smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ via the following diagram:
@@
This point of view does not affect the wavefront set analysis. As a reminder, the wavefront set $\text{WF}(u)$ of a distribution $u\in\mathcal{D}^{\prime}(SM)$ describes the directions in $T^{*}(SM)$ along which the distribution is not smooth. It refines the notion of the singular support of $u$, which only records the points in $SM$ where $u$ is singular. For a detailed introduction, we refer the reader to [Hör03, Chapter 8].
@@
Let $q\in\mathcal{C}^{\infty}(\widetilde{S}M,\mathbb{R}_{>0})$ be the function such that $\phi^{*}\omega=q\tilde{\omega}$. Then, we get
$\Phi u$ $=\widetilde{L}^{-1}\phi^{*}Lu$
$=\widetilde{L}^{-1}\phi^{*}(u\omega)$
$=\widetilde{L}^{-1}(\phi^{*}(u)q\tilde{\omega})$
@@
Since multiplication by the nowhere-vanishing function $q$ is elliptic, we get the result by elliptic regularity (see [Hör03, Theorem 8.3.2]). ∎
By the properties of wavefront sets under pullback operators (see [Hör03, Theorem 8.2.4] for instance), we thus obtain
$\textup{WF}(\Phi u)=d\phi^{\top}(\textup{WF}(u))$
for all $u\in\mathcal{D}^{\prime}_{\textup{tr}}(SM)$, where $d\phi^{\top}:T^{*}(SM)\to T^{*}(\widetilde{S}M)$ is the symplectic lift of $\phi^{-1}$ to the cotangent bundles given by
$d\phi^{\top}(v,\xi)\coloneqq\left(\phi^{-1}(v),d\phi^{\top}_{\phi^{-1}(v)}\xi\right),\qquad(v,\xi)\in T^{*}(SM).$
2.1.4. Vertical Fourier expansions
@@
$\mathcal{C}^{\infty}(SM)=\oplus_{k\in\mathbb{Z}}\Omega_{k},\qquad\Omega_{k}\coloneqq\left\{u\in\mathcal{C}^{\infty}(SM)\mid\,Vu=iku\right\}.$
This decomposition is orthogonal with respect to the $L^{2}$ inner product on $SM$ and with $\mathcal{C}^{\infty}$ being replaced by $L^{2}$. For any $u\in\mathcal{C}^{\infty}(SM)$, we shall write $u=\sum_{k\in\mathbb{Z}}u_{k}$, where the $k$th Fourier mode $u_{k}\in\Omega_{k}$ is given by
@@
$\langle u_{k},\varphi\rangle\coloneqq\langle u,\varphi_{-k}\rangle_{\mathcal{D}^{\prime}(SM)},\qquad\varphi\in\mathcal{C}^{\infty}(SM)$
@@
It also worth noting that the ladder operators $\eta_{\pm}$ in equation (2.1) take their name from the fact that they act as raising/lowering operators on the Fourier decomposition, that is,